# A Systems Thinking Approach to Algorithmic Fairness

### v2

@@ 4. Modeling fairness as a complex system
Application Protected Attribute $A$ Mediator $W$ Data $X$ Decision $D$ Outcome $Y$
Credit Race, gender Creditworthiness Income, credit history Deny loan? Loan default?
Insurance Race, gender Insurability Demographics, claims history Deny policy? Claim filing?
Housing Race, gender Home value Location, size, condition Home valuation? Final sales price?
Employment Race, gender Qualifications Experience, education Screen resume? Turnover?
Education Race, gender Merit Grades, test scores Reject applicant? Graduation?
Research Race, gender Merit Quality, clarity, originality Reject paper? No. of citations?
@@
We begin with Figure 1a, which is a block diagram representation of supervised machine learning. In the center of this block diagram is a black box machine learning model that receives data $X$ as input and returns decision $D$ as output. The decision $D$ leads to some outcome $Y$, which becomes future data $X$ through a feedback loop. A key question is whether a protected attribute $A$ should or should not be fed into the black box model and if so, then how.
Next up we have Figure 1b, which is a causal Bayesian network representation (CBN) of causal inference. CBNs were pioneered by Pearl(Pearl, 2000; Pearl and Mackenzie, 2018; Pearl et al., 2016) and allow us to symbolically encode knowledge and assumptions about causal effects between variables, which are shown in this graph as nodes. The first node of the CBN represents a protected attribute $A$, which has a direct effect on a mediator $W$ ($A\rightarrow W$). The mediator $W$ has a direct effect on an outcome $Y$ ($W\rightarrow Y$). We assume that the mediator $W$ completely explains away the causal effect between the protected attribute $A$ and the outcome $Y$ ($Y\perp\!\!\!\perp A|W$). However, the mediator $W$ might be considered biased by some people if they attribute the disparities across groups as being external. The mediator $W$ is filled white because it is latent, but it can still be partially measured as data $X$ ($W\rightarrow X$). Data $X$ is then fed into a machine learning model that is embedded inside the node for decision $D$ ($X\rightarrow D$). Finally, the decision $D$ also has a direct effect on an outcome $Y$ ($D\rightarrow Y$). CBNs are represented by directed acyclic graphs (DAGs), which means that we cannot use them to model feedback loops.
Finally we have Figure 1c, which is a causal loop diagram (CLD) representation of system dynamics. CLDs were pioneered by Forrester (Forrester, 1971) and are similar to CBNs in that they allow us to symbolically encode knowledge and assumptions about cause and effect relationships between variables(Sterman, 2000). But unlike CBNs, CLDs allow us to model feedback loops. This is useful for modeling why we might believe that disparities exist between protected groups. In the center of the CLD is a node for decision $D$, which like before contains our machine learning model. This CLD is two sided, with the left side representing group A and the right side representing group B. Data from group A $X_{A}$ and group B $X_{B}$ is fed into the ML model, which renders a decision $D$ ($X_{A}\rightarrow D\leftarrow X_{B}$). In this scenario, a decision $D$ that leads to a more favorable outcome for group A $Y_{A}$ might also lead to a less favorable outcome for group B $Y_{B}$ ($Y_{A}\leftarrow D\rightarrow Y_{B}$). This could happen on an absolute basis if both groups are competing for the same resources (e.g. a seat at an elite university). This could also happen on a relative basis if one group receives proportionately more favorable decisions than another (e.g. one group receives higher loan approval rates than another). Note that there is a time delay (denoted by ——) between decision $D$ and outcome $Y$. The outcome $Y$ is recorded as future data $X$ ($Y_{A}\rightarrow X_{A}$ and $Y_{B}\rightarrow X_{B}$), which in turn drives a future decision $D$ ($X_{A}\rightarrow D\leftarrow X_{B}$). We represent this using two reinforcing feedback loops (R1 and R2), which we will show later results in a recurring pattern of behavior called the “success to the successful” archetype.
Let’s say that group A has historically been more successful than group B, which was due to some unjust practice from the past. Then when a model is making a decision $D$ about how to allocate resources to individuals based on past data $X_{A}$ and $X_{B}$ ($X_{A}\rightarrow D\leftarrow X_{B}$), it may allocate more resources to individuals from group A since it would be using that group’s past success to predict future success. This could breed resentment among members of group B, who might demand the use of their protected attribute $A_{B}$ to intervene on a decision $D$ ($A_{B}\rightarrow D$). Therefore, rather than the ML model making a decision $D$ on data $X$ alone, it might use the protected attribute $A_{B}$ to affirmatively favor group B. However, this affirmative action policy might have an unintended consequence that could harm group B’s future outcomes $Y_{B}$ ($A\rightarrow Y_{B}$), offsetting the benefits of an affirmative action policy. We represent this for both sides using a series of balancing feedback loops (B1 and B2) and reinforcing feedback loops (R3 and R4), which we will show later results in a recurring pattern of behavior called the “shifting the burden / addiction” archetype.
These causal graphs form the basis for modeling the complex political and legal debate over how society can build fair machine learning models. They allow us to model the social structures that influence human behavior over time. In the next several sections, we will dive deeper into how to model bias and discrimination using these causal graphs. They will allow us to understand the complex tradeoffs that exist from different fairness policies, providing a set of mathematical rules for policymakers on how to design AI policies to achieve their political goals.
5. Modeling fairness as a causal Bayesian network
Causal Bayesian networks (CBNs) can be used as a higher level representation of any supervised machine learning model. It is at this layer of abstraction that we can begin modeling the different biases that society might have. By understanding these various sources of bias, we can model the different ways that a model can discriminate. Finally, we can use these insights to help policymakers on both sides of the aisle to design AI policies for machine learning models that they would deem to be fair.
@@ 5. Modeling fairness as a causal Bayesian network
But if we instead believe that the data is fundamentally biased (external attribution), then fairness through unawareness could lead to covert discrimination (Figure 2d). In this scenario, there is bias between the protected attribute $A$ and the mediator $W$ ($A\rightarrow W$), which flows across the entire graph and causes all of the nodes in the graph to be colored red. This bias can be attributed to structural or systemic forms of discrimination that unfairly disadvantages one group over another. For example, if a disadvantaged group was overtly discriminated against in the past, then the present data $X$ may still reflect that negative bias. By removing the edge from the protected attribute $A$ to the decision $D$, it could reinforce one group’s prior success over another. Covert discrimination could form the basis of a disparate impact or indirect discrimination claim. We would say that the protected attribute $A$ has a negative indirect effect on the decision $D$ ($A\rightarrow W\rightarrow X\rightarrow D$), leading to harmful discrimination.
@@
But if we disagree with the belief that the data is fundamentally biased (internal attribution), then fairness through affirmative action could be perceived as a form of reverse discrimination (Figure 2f). The intervention of the protected attribute $A$ on the decision $D$ ($A\rightarrow D$) could lead to individuals receiving a positive decision $D$ but a negative outcome $Y$. For example, a machine learning model may predict that an individual from a historically disadvantaged group may end up with a negative outcome $Y$. However, this would lead to a greater disparity in decisions across protected groups. So the model may instead use knowledge of that individual’s membership in a protected group $A$ to render a more positive decision $D$. But this positive decision $D$ could lead to a worse outcome $Y$ for the individual than if that individual had just received a negative decision $D$ in the first place. An example might be approving a minority borrower for a loan even if that borrower is not creditworthy, which could result in a loan default. We would say that the protected attribute $A$ has a positive direct effect on the decision $D$ ($A\rightarrow D$), leading to harmful discrimination.
@@
We now move on to the more extreme forms of fairness. Reverse discrimination could lead to fairness through supremacism (Figure 2a). An advantaged group may use this approach to actively suppress a disadvantaged group (e.g. white supremacism). This could happen if an advantaged group believes that a disadvantaged group has been unfairly receiving preferential treatment. For example, if a disadvantaged group has benefited from reverse discrimination in the past, then the present data $X$ may still reflect that positive bias and is colored blue. The protected attribute $A$ is then used to correct for that positive bias in the decision $D$ ($A\rightarrow D$) and is colored red. This could explain what happened when hiring managers racially discriminated against job applicants with Black sounding names compared to those with White sounding names even through the resumes were the same, as shown by Bertrand and Mullainathan(Bertrand and Mullainathan, 2003). We would say that the protected attribute $A$ has a positive indirect effect on the decision $D$ ($A\rightarrow W\rightarrow X\rightarrow D$) and a negative direct effect on the decision $D$ ($A\rightarrow D$), thus canceling out the biases and making the decision fair.
@@
Another extreme form of fairness is fairness through lottery (Figure 2g). If we believe that the data is fundamentally biased, an alternative approach to removing bias in the decision $D$ is to remove the edge from data $X$ to decision $D$. The advantage of this approach to fairness is that it provides stronger guarantees of equalizing outcomes across groups. This is particularly effective for dealing with intersectionality issues, ensuring that every potential intersection of a protected class (e.g. race, gender, and religion) receives the same decision rates. Individuals cannot change the decision by altering their own data $X$. The decision $D$ could be assigned randomly or there could be a fixed rule where everyone receives a positive decision $D$ or a negative decision $D$. We would say that the protected attribute $A$ has no direct effect on the decision $D$ and its indirect effects have been canceled out, making the decision fair.

### v3

@@ 5. Modeling fairness as a causal inference problem
Application Mediator $W$ Data $X$ Decision $D$ Outcome $Y$
Credit Creditworthiness Income, credit history Deny loan? Loan default?
Insurance Insurability Demographics, claims history Deny policy? Claim filing?
Housing Home value Location, size, condition Home valuation? Final sales price?
Employment Qualifications Experience, education Screen resume? Turnover?
Education Merit Grades, test scores Reject applicant? Graduation?
Research Merit Quality, clarity, originality Reject paper? No. of citations?
@@
Figure 3. Modeling bias, fairness, and discrimination using causal Bayesian networks (CBNs)
@@
But if we instead believe that the data is fundamentally biased (external attribution), then fairness through unawareness could lead to covert discrimination (Figure 3d). In this scenario, there is bias between the protected attribute $A$ and the mediator $W$ ($A\rightarrow W$), which flows across the entire graph and causes all of the remaining nodes in the graph to be colored red. This bias can be attributed to structural or systemic forms of discrimination that unfairly disadvantages one group over another. For example, if a disadvantaged group was overtly discriminated against in the past, then the present data $X$ may still reflect that negative bias. By removing the edge from the protected attribute $A$ to the decision $D$, it could reinforce one group’s prior success over another. Covert discrimination could form the basis of a disparate impact or indirect discrimination claim. We would say that the protected attribute $A$ has a negative indirect effect on the decision $D$ ($A\rightarrow W\rightarrow X\rightarrow D$), leading to harmful discrimination.
@@
But if we disagree with the belief that the data is fundamentally biased (internal attribution), then fairness through affirmative action could be perceived as a form of reverse discrimination (Figure 3f). The intervention of the protected attribute $A$ on the decision $D$ ($A\rightarrow D$) could lead to individuals receiving a more positive decision $D$ but a more negative outcome $Y$. For example, a machine learning model may predict that an individual from a historically disadvantaged group may end up with a negative outcome $Y$. However, this would lead to a greater disparity in decisions across protected groups. So the model may instead use knowledge of that individual’s membership in a protected group $A$ to render a more positive decision $D$. But this positive decision $D$ could lead to a worse outcome $Y$ for the individual than if that individual had just received a negative decision $D$ in the first place. An example might be approving a minority borrower for a loan even if that borrower is not creditworthy, which could result in a loan default. We would say that the protected attribute $A$ has a positive direct effect on the decision $D$ ($A\rightarrow D$), leading to harmful discrimination.
@@
We now move on to the more extreme forms of fairness. Reverse discrimination could lead to fairness through supremacism (Figure 3a). An advantaged group may use this approach to actively suppress a disadvantaged group (e.g. white supremacism). This could happen if an advantaged group believes that a disadvantaged group has been unfairly receiving preferential treatment. For example, if a disadvantaged group has benefited from reverse discrimination in the past, then the present data $X$ may still reflect that positive bias and is colored blue. The protected attribute $A$ is then used to correct for that positive bias in the decision $D$ ($A\rightarrow D$) and is colored red. This might explain what happened when hiring managers racially discriminated against job applicants with Black sounding names compared to those with White sounding names even through the resumes were the same, as shown by Bertrand and Mullainathan(Bertrand and Mullainathan, 2003). We would say that the protected attribute $A$ has a positive indirect effect on the decision $D$ ($A\rightarrow W\rightarrow X\rightarrow D$) and a negative direct effect on the decision $D$ ($A\rightarrow D$), thus canceling out the biases and making the decision fair.
@@
Another extreme form of fairness is fairness through lottery (Figure 3g). If we believe that the data is fundamentally biased, an alternative approach to removing bias in the decision $D$ is to remove the edge from data $X$ to decision $D$. The advantage of this approach to fairness is that it provides stronger guarantees of equalizing decision rates across groups. This is particularly effective for dealing with intersectionality issues, ensuring that every potential intersection of a protected class (e.g. race, gender, and religion) receives the same decision rates. Individuals cannot change the decision by altering their own data $X$. The decision $D$ could be assigned randomly or there could be a fixed rule where everyone receives a positive decision $D$ or a negative decision $D$. We would say that the protected attribute $A$ has no direct effect on the decision $D$ and its indirect effects have been canceled out, making the decision fair.
@@
5.1. Mapping fairness to politics
Table 2. Mapping fairness type to the political spectrum
Political worldview Far- left Left- wing Right- wing Far- right
Type Lottery Affirmative action Unaware- ness Suprem- acism
Individual fairness Yes No Yes No
Group fairness Yes Yes No No
Pearl’s law First Second
[…]
Right-wing notions of fairness rely on fairness through unawareness, which is associated with the criterion $D\perp\!\!\!\perp A|W$. That is, the decision $D$ is conditionally independent of the protected attribute $A$ given some mediator $W$. For example, a college admissions decision is independent of race given an applicant’s merit. This leads to individual fairness, where similar individuals are treated similarly regardless of protected group membership(Castelnovo et al., 2022; Verma and Rubin, 2018; Mehrabi et al., 2021; Binns, 2019). This form of fairness relies on Pearl’s second fundamental law of causal inference: conditional independence.
Left-wing notions of fairness rely on affirmative action, which is associated with the criterion $D\perp\!\!\!\perp A$. That is, the decision $D$ is independent of the protected attribute $A$, which is achieved by using the protected attribute $A$ to correct for biases in the data $X$ that is used in the decision $D$. We imagine what the applicant’s decision would have been had they been a member of a counterfactual class. We then perform an intervention to obtain the same decision. For example, a college admissions decision would affirmatively use race to correct for deficiencies in the applicant’s merit due to structural discrimination. This leads to group fairness, where different groups receive more similar aggregate decision rates(Castelnovo et al., 2022; Verma and Rubin, 2018; Mehrabi et al., 2021; Binns, 2019). This form of fairness relies on Pearl’s first fundamental law of causal inference: counterfactuals and interventions.
5.2. Mapping discrimination to antidiscrimination law
Table 3. Mapping discrimination type to antidiscrimination law
US Type Disparate treatment Disparate impact
EU Type Direct discrimination Indirect discrimination
Overt X
Covert X
Reverse X
“No” X
[…]
A table describing how to map causal forms of discrimination with legal forms of discrimination.
When a protected attribute $A$ has an indirect effect on a decision $D$, we might consider this as a form of disparate impact or indirect discrimination. This would map to covert discrimination (Figure 3d) and “no” discrimination (Figure 3h). Even though the protected attribute $A$ has no direct effect on a decision $D$ (making the policy neutral), the inclusion of biased data or exclusion of unbiased data could cause the protected attribute $A$ to have an indirect effect on the decision $D$ that adversely affects a protected group.
For disparate impact in the US, there is a three-step test(of Justice Civil Rights Division, 2024). The first step is to establish that a neutral policy has an adverse disparate impact on a protected group. This would involve testing whether there is a disparity in decisions $D$ across the protected class $A$. The second step is to establish the business necessity of that policy. This would involve arguing that the data $X$ being used has some causal relationship to the mediator $W$. The final step is to identify a less discriminatory alternative model, which would involve performing a model search.
CBNs provide an analytical framework to reason about whether a model is causing illegal discrimination. By being able to explicitly model the different forms of discrimination, CBNs provide a common language that businesses and regulators can use to ensure that machine learning models are compliant with laws and regulations. Appendix B describes the normative effects of the CBN from the different forms of fairness and discrimination.
5.3. The limitations of causal Bayesian networks
CBNs have a key limitation: they can only be represented using directed acyclic graphs (DAGs). Recall from the block diagram in Figure 1 that outcome $Y$ could be recorded as future data $X$. We can modify the CBN from Figure 2 to model this feedback loop by adding a edge from outcome $Y$ to the mediator $W$ as shown in Figure 4 (this edge has been colored red), which is then recorded as data $X$. Even though this violates the DAG constraint of CBNs, we still need to model how different fairness policies may perpetuate disparities across protected groups over time. In the next section, we will show how to overcome this limitation using causal loop diagrams.
[…]
Figure 4. Adding a feedback loop to the causal Bayesian network
[…]
A causal Bayesian network where the outcome $Y$ causes mediator $W$, creating a directed acyclic graph violation.
6. Modeling fairness as a system dynamics problem
We can transform the block diagram from Figure 1 and the CBN from Figure 2 into a causal loop diagram (CLD) as shown in Figure 5. CLDs were pioneered by Forrester (Forrester, 1971) and are similar to CBNs in that they allow us to symbolically encode knowledge and assumptions about cause and effect relationships between variables(Sterman, 2000). But unlike CBNs, CLDs allow us to model feedback loops. This is useful for modeling why we might believe that disparities continue to exist between protected groups.
[…]
Figure 5. Simplified causal loop diagram representation of algorithmic fairness
[…]
A causal loop diagram showing a decision $D$ being caused by data $X$ and the protected attribute $A$.
The CLD in Figure 5 is a simplified version of the full CLD that we will present later in this section. This shows how the same variables in the block diagram and the CBN can be mapped onto a CLD. In the center of the CLD is a node for decision $D$, which like before contains our machine learning model. This CLD is two sided, with the left side representing group A and the right side representing group B. Data from group A $X_{A}$ and group B $X_{B}$ is fed into the ML model, which renders a decision $D$ ($X_{A}\rightarrow D\leftarrow X_{B}$). Note that there is a time delay (denoted by ——) between decision $D$ and outcomes $Y_{A}$ and $Y_{B}$. The outcomes $Y_{A}$ and $Y_{B}$ are recorded as future data $X_{A}$ and $X_{B}$ ($Y_{A}\rightarrow X_{A}$ and $Y_{B}\rightarrow X_{B}$), which in turn drives a future decision $D$ ($X_{A}\rightarrow D\leftarrow X_{B}$). The outcomes $Y_{A}$ and $Y_{B}$ also influence whether the protected attribute $A_{A}$ or $A_{B}$ might be used to influence a decision $D$. For example, if group B has historically received worse outcomes that group A, then the group might demand the protected attribute $A_{B}$ to be affirmatively used to modify decision $D$ (fairness through affirmative action). However, the use of the protected attribute $A_{A}$ or $A_{B}$ could also impact outcomes $Y_{A}$ or $Y_{B}$.
Within a CLD, every path is labeled as having a positive (+) or negative (–) direction. If the path is positive, then changes in one variable leads to changes in another variable in the same direction. For example, an increase in one variable would lead to an increase in the other variable. Alternatively, a decrease in one variable would lead to a decrease in the other variable. If the path is negative, then changes in one variable leads to changes in another variable in the opposite direction. For example, an increase in one variable would lead to a decrease in the other variable. Alternatively, a decrease in one variable would lead to an increase in the other variable.
Figure 6. Success to the successful archetype
[…]
Modeling fairness using the success to the successful archetype.
[…]
Figure 7. Limits to success archetype
[…]
Modeling fairness using the limits to success archetype.
Figure 8. Shifting the burden / addiction archetype
[…]
Modeling fairness using the success to the successful archetype.
In Figure 9, we combine these three archetypes into a single CLD to model the algorithmic fairness problem. This diagram is an expanded version of the CLD in Figure 5c. We begin with the “success to the successful” archetype. A machine learning model is embedded inside decision $D$, which makes allocation decisions based on decision $X$ and has a tendency to favor one group over another. Those decisions lead to resources being assigned to each group, which eventually determines that group’s success and is measured as an outcome $Y$. Past success $Y$ becomes future data $X$, which is then fed back into the machine learning model to drive a future allocation decision $D$. This is represented using reinforcing loops R1 and R2. Since the allocation decisions $D$ are initially only driven by data $X$, success to the successful is equivalent to fairness through unawareness (Figure 3c).
Allocation to Ainstead of B ($D$)Resourcesto AResourcesto BSuccessof A ($Y_{A}$)Successof B ($Y_{B}$)Data fromA ($X_{A}$)Data fromB ($X_{B}$)R1R2++++++Limitingaction on ALimitingaction on BB1B2++Constrainton AConstrainton BExternalinterventionby A ($A_{A}$)Externalinterventionby B ($A_{B}$)+B3B4Dependenceon externalinterventionby ADependenceon externalinterventionby B++R3R4MLModel————
[…]
Finally, we add in the “shifting the burden / addiction” archetype. If one group is allocated less resources than another group, then it could breed resentment by the disadvantaged group. That group could then demand that the protected attribute $A$ be used to externally intervene on the allocation decision $D$. This external intervention is equivalent to fairness through affirmative action (Figure 3e). While it might reduce allocation disparities in the short term, it may actually create a dependence on the external intervention over the long term. This is because if allocation decisions $D$ are not based purely on data $X$, which is a measure of mediator $W$ (e.g. merit, qualifications), then the assignment of resources to the disadvantaged group may not lead to their success $Y$. For example, approving a loan to a minority applicant even though the applicant is not creditworthy could lead to a worse outcome (default) than if the loan was denied in the first place. That decreased success in outcome $Y$ for the less successful group could trigger a demand for an even stronger external intervention $A$ on the decision $D$, reinforcing this problem. In the meanwhile, attention may be taken away from the internal solution of optimizing for the mediator $W$ and the data $X$ so that the disadvantaged group would perform better under fairness through unawareness. This may especially be the case if the disadvantaged group believes that the mediator $W$ is biased and therefore does not focus on optimizing that metric. Appendix C describes the positive effects of the CLD from the different forms of fairness and discrimination.
7. Modeling fairness as a causal hierarchy
Layer 1. Association
Activity:
Seeing, observing
Layer 2. Intervention
Activity:
Doing, intervening
Layer 3. Counterfactuals
Activity:
Imagining, retrospection, understanding
Layer 1. Association
Field:
Supervised machine learning
Representation: Block diagrams Application: Prediction
Layer 2. Acyclic causation
Field:
Causal inference (Judea Pearl)
Representation: Causal Bayesian networks Application: Control
Layer 3. Cyclic causation
Field:
System dynamics (Jay Forrester)
Representation: Causal loop diagrams Application: Simulation
[…]
Figure 10. Causal hierarchy representation of algorithmic fairness
[…]
Adding a system dynamics layer to Pearl’s causal hierarchy.
Each of the layers in this new causal hierarchy have unique properties and behaviors that make them optimized for different applications. The first layer is associated with supervised machine learning and is optimized for prediction. The second layer is associated with causal inference and is optimized for control. The third layer is associated with system dynamics and is optimized for simulation. These three new causal layers make up an expanded causal hierarchy that is consistent with systems thinking principles. Appendix D shows how this can be part of a larger systems map.
8. Conclusion
Systems thinking provides us with a novel way to model the algorithmic fairness problem by “thinking outside the black box.” While a large amount of AI research has been focused on the hidden layers that are “inside” a black box model, this paper has focused on the causal layers that are “outside” a black box model. These outer layers are where we can encode our knowledge and assumptions about where we might believe that bias occurs in the data generating process. We showed that the first outer layer can be represented using causal Bayesian networks and the second outer layer can be represented using causal loop diagrams. This gives us an elegant way of modeling the emergent behavior between machine learning, causal inference, and system dynamics.
This approach allow us to develop a more unified and holistic view of the fairness and machine learning problem, providing essential connections between computer science, politics, and the law. This is the key to building AI systems that are aligned with society’s democratic values and the policymakers’ political agendas. It will enable businesses and regulators to communicate with each other using a common causal language, one that can be used by both data scientists and lawyers. These models may even aid in AI deregulation by identifying mathematical flaws in the existing laws and regulations. This will help businesses to comply with antidiscrimination laws and regulations, reducing business uncertainty and expensive litigation while providing a much needed layer of transparency and trust for consumers of AI products and services.
Systems thinking supplies the missing mathematical models that are needed to build responsible AI systems. To solve fairness, we need to bring together both the social and technical aspects of what is ultimately a complex systems problem. On the social side, it requires modeling the diversity of perspectives that come from both sides of the political spectrum. On the technical side, it requires unifying analytical techniques that were previously thought of as being independent. By unifying all of these concepts together, we hope that this enables the development of AI policies and systems that could lead to practical, real-world solutions for reducing disparities across protected groups in society.
Appendix A Case study: Racial disparities in the United States
We can use the causal models presented in the main part of this paper to demonstrate how to model the complex political debate over racial disparities in the US. According to the US Census Bureau, median household income in 2021 was \$100,573 for Asians, \$74,932 for Whites, \$57,671 for Hispanics, and \$48,297 for Blacks(Survey, 2021). Meanwhile, median household wealth in 2021 was \$320,900 for Asians, \$250,400 for Whites, \$48,700 for Hispanics, and \$27,100 for Blacks(Kochnar and Moslimani, 2023). We would like to understand why Whites and Asians have higher median household income and wealth than Blacks and Hispanics.
When it comes to understanding the root causes of racial disparities, we can take into account some of the original sources of bias in the data. We know that the US has historically had unjust policies that overtly discriminated against non-Whites. These included overtly racist policies like slavery, Jim Crow segregation, and the Chinese Exclusion Act. For convenience, we replicate the CBN for overt discrimination in Figure 11.
[…]
Figure 11. Overt discrimination
[…]
Figure 12. Feedback from overt discrimination
[…]
A CBN with a feedback loop from the outcome $Y$ to the mediator $W$.
[…]
Figure 13. Covert discrimination
[…]
This creates a dilemma for policymakers: Should the US institute fairness through unawareness, even if such a policy could replicate past racial disparities (i.e. the “success to the successful” archetype as shown in Figure 6)? Or should the US institute affirmative action to try to correct for past injustices (i.e. the “shifting the burden” archetype as shown in Figure 8)?
Due to the “success to the successful” archetype, those racial groups that have traditionally been left behind may not be able to compete against other racial groups that are presently more successful. For example, Whites and Asians have access to better funded public schools, safer neighborhoods, higher quality healthcare, and better social networks for job opportunities. These advantages could result in Whites and Asians being able to outcompete Blacks and Hispanics for the same resources under fairness through unawareness, since there will be a disparity in their merit and qualifications.
So let’s say that US policymakers (particularly Democrats) succeed in instituting affirmative action as shown in Figure 14. While affirmative action might help correct for biases in the decision $D$, notice in Figure 14 that it does not necessarily correct for biases in the outcome $Y$. The protected attribute $A$ does not have a direct effect on the outcome $Y$. It can only directly affect the decision $D$. Correcting for biases in the decision $D$ cannot fully offset biases in the mediator $W$, which means that the outcome $Y$ remains biased. The path $D\rightarrow Y$ only has a weak effect on the outcome $Y$. On the other hand, the path $W\rightarrow Y$ has a strong effect on the outcome $Y$. This is shown in Figure 15.
[…]
Figure 14. Fairness through affirmative action
[…]
Figure 15. Strong and weak effects
[…]
For example, an affirmative action policy in college admissions might help more Blacks and Hispanics to be accepted ($A\rightarrow D$). But because they were not accepted purely based on their merit (mediator $W$), they might be less likely to graduate (outcome $Y$). This is because they might be put into a class where their White and Asian counterparts would consistently outperform them. Another example could be an affirmative action policy for credit decisions. This might help more Blacks and Hispanics to receive more loans ($A\rightarrow D$). But since they were not accepted purely based on their creditworthiness (mediator $W$), they would be more likely to default (outcome $Y$). This is because affirmative action does not directly improve one’s merit or creditworthiness. Even worse, affirmative action could have the opposite effect on the mediator $W$.
[…]
Figure 16. Internal solution and external intervention
[…]
Placing Blacks and Hispanics into higher level positions than is warranted by their merits and qualifications could also exact a harmful psychological toll. This could reinforce their views that they are victims of an unjust system(Kaufman, 2020), leading to a sense of learned helplessness(Seligman, 1972) and a loss of agency. This could result in a reduction of human capital and productivity, which is the primary determinant of income and wealth in the primarily capitalist economic system of the US(Becker, 1964). Rather than helping Blacks and Hispanics become more competitive against Whites and Asians, affirmative action may have instead made them less competitive.
Asians in the US have not historically benefited from affirmative action policies. Asians have instead relied on fairness through unawareness, as shown in Figure 17. In some cases like college admissions, they may even have to outperform their White counterparts to receive the same admission decisions(Grossman et al., 2023). Asians as a group have tended to focus their efforts on the internal solution of competing based on their merits and qualifications. By working against Asians, affirmative action programs may have actually caused Asians to work even harder than other racial groups. By focusing on developing their own human capital instead of demanding strong government intervention, not only have Asians closed the gap in income and wealth with Whites over the long-term, they have actually surpassed them. Over time, this has also reduced perceived biases in the data between Asians and Whites.
[…]
Figure 17. Fairness through unawareness
[…]
Due to the “limits to success” archetype, human capital represents a constraint on the level of success for any racial group. By focusing more on education and pursuing higher income careers (e.g. STEM) than other racial groups, Asians have been able to catch up and surpass Whites in terms of income and wealth. This may reflect a selection bias in the types of Asians who emigrate to the US, who oftentimes come to pursue better educational and career opportunities.
This is the reason why some US policymakers (particularly Republicans) have been arguing in favor of fairness through unawareness. While acknowledging that historically unjust policies like slavery and Jim Crow have created the initial disparities between Whites and non-Whites, they also recognize that external government interventions like affirmative action may cause more harm than good. They would argue that it is not an accident that the racial group that has benefited the most from affirmative action (i.e. Blacks) ultimately received the worst outcomes. The focus by the Black community on how past injustices perpetuate racial disparities may create a self-fulfilling prophecy. Meanwhile, the racial group that has benefited the least from affirmative action (i.e. Asians) ultimately received the best outcomes. The focus by the Asian community on developing their own human capital through education and hard work (i.e. the model minority) has allowed that community to close their initial disparities with Whites.
On the other hand, some US policymakers (particularly Democrats) have argued for a doubling down on affirmative action and DEI programs. This was especially the case during the past four years under the Biden administration. To address the fact that affirmative action only has a weak effect on outcomes, their goal was to deploy affirmative action and DEI programs across the entire US economy. This would create a racially segregated economic system that provides special treatment for certain minority groups. By instituting a system of double standards and quotas across the US economy, their hope was to equalize outcomes across the racial groups. Doing so would have required a radical restructuring of the US economy, from one that primarily relies on capitalism to one that primarily relies on segregation. It would have reduced the size of the overall economic pie as opportunities are distributed not primarily based on qualifications but instead to reach diversity quotas. As we saw in the recent election, there was not enough popular support in the US for such a radical transformation of the US economy.
Affirmative action is a low-leverage solution that at best can only modestly decrease racial disparities. It is not a fundamental, high-leverage solution for reducing racial disparities over the long-term. As long as there are racial disparities in the mediator $W$, then there will be racial disparities in the outcome $Y$. This would mean that racial disparities in household income and wealth in the US would continue to persist. The focus should instead be on closing disparities in the mediator $W$. For example, Democrats could call for increased government funding of education, training, and health programs to help minorities improve their qualifications so that they can be more productive. On the other hand, Republicans would continue their calls for the banning of racial preferences in high-stakes decisions to keep all racial groups accountable to the same high standards.
These models provide a basis for acknowledging that while historical policies like slavery and Jim Crow generated the initial racial disparities between Whites and Blacks, contemporary policies like affirmative action and DEI may have reinforced those disparities. Affirmative action should at most be used as a temporary solution for marginally closing the racial equity gap. If it is used, then there should be a long-term plan to eventually switch back to fairness through unawareness. If an addictive dependency on affirmative action develops that diverts attention away from internal solutions for closing the merit and qualifications gap, then the affirmative action policy should be ended. Systems thinking shows us that AI researchers should not dismiss fairness through unawareness as a fundamental solution for closing the racial equity gap.
Appendix B Normative effects of fairness and discrimination